FIND THE MEASURES OF 2 SUPPLEMENTARY ANGLES IF THE MEASURE OF ONE ANGLE IS 5 LESS THAN 4 TIMES THE MEASURE OF THE OTHER.
Answer by mananth(16946)

let one angle be x deg
the other will be 4x-5
x+4x-5=180
5x-5=180
5x=185
x=37
the other angle = 180-37= 143deg